# via Object-Centric Visual Representation Learning

### v2

@@ Abstract
Video Scene Graph Generation (VidSGG) aims to represent dynamic visual content by detecting objects and modeling their temporal interactions as structured graphs. Prior studies typically target either coarse-grained box-level or fine-grained panoptic pixel-level VidSGG, often requiring task-specific architectures and multi-stage training pipelines. In this paper, we present UNO (UNified Object-centric VidSGG), a single-stage, unified framework that jointly addresses both tasks within an end-to-end architecture. UNO is designed to minimize task-specific modifications and maximize parameter sharing, enabling generalization across different levels of visual granularity. The core of UNO is an extended slot attention mechanism that decomposes visual features into object and relation slots. To ensure robust temporal modeling, we introduce object temporal consistency learning, which enforces consistent object representations across frames without relying on explicit tracking modules. Additionally, a dynamic triplet prediction module links relation slots to corresponding object pairs, capturing evolving interactions over time. We evaluate UNO on standard box-level and pixel-level VidSGG benchmarks. Results demonstrate that UNO not only achieves competitive performance across both tasks but also offers improved efficiency through a unified, object-centric design. Our code will be fully released upon acceptance.
@@ 1 Introduction
Video Scene Graph Generation (VidSGG) aims to extract structured, dynamic representations from videos by modeling objects as nodes and their pairwise interactions as edges in spatio-temporal graphs. These structured representations offer both interpretability and compositionality, making VidSGG a critical component in various downstream tasks such as video understanding [17, 49, 58, 41], video reasoning [62, 36] and robotic reasoning [55].
@@
Current VidSGG research primarily follows two directions, distinguished by the level of visual granularity: box-level VidSGG (also referred to as Dynamic Scene Graph Generation or DSGG) [17, 47, 32, 25, 45], and panoptic pixel-level VidSGG (also known as Panoptic Video Scene Graph Generation or PVSG) [58, 50, 57]. The former focuses on coarse-grained object representations using bounding boxes and typically models relationships at the frame level. The latter provides fine-grained, pixel-level representations using panoptic segmentation masks, where object trajectories are treated as graph nodes and interactions—including object-object and object-background—are captured throughout the video. Importantly, PVSG emphasizes the temporal consistency of object identities across frames.
@@
In scenarios requiring multi-level scene understanding, a unified VidSGG model capable of handling both tasks is highly desirable. Such a model could flexibly adapt to diverse visual representations and support a wider range of applications without task-specific architectural redesign. However, achieving this unification is non-trivial due to the differing structural assumptions, temporal modeling requirements, and visual encoding strategies inherent to each task. Prior attempts have relied on multi-stage pipelines involving either box-level or pixel-level representation, followed by a tracking module [25, 48, 59, 58, 34], which introduce significant computational overhead and often result in suboptimal performance due to decoupled learning and limited parameter sharing. On the other hand, designing a unified, end-to-end solution encounters the core challenge of learning a semantically consistent spatio-temporal representation that generalizes across varying levels of granularity while remaining aligned with the underlying video dynamics.
@@
Beyond spatial granularity, VidSGG approaches also differ in their temporal-level representations, which can be broadly categorized into frame-level [7, 25, 32, 47] and tracklet-level [58, 34] methods. Frame-level methods construct scene graphs independently at each frame, aligning naturally with box-level VidSGG, but often fall short in modeling long-term interactions and maintaining temporal consistency. In contrast, tracklet-level methods are more accurate because they link object instances across frames to capture temporal dynamics explicitly, a strategy more common in pixel-level VidSGG. This lack of temporal generality hinders their ability to capture coherent and continuous scene dynamics—especially in applications requiring both fine-grained relationship modeling and long-range temporal reasoning. A detailed comparison of UNO with prior works is presented in Tab. 1.
To address this challenge, we propose UNO, a unified, object-centric, single-stage VidSGG framework that effectively supports both box-level and pixel-level tasks. Fig. 1 illustrates the overall concept of UNO. Our central hypothesis is that despite their differences, DSGG and PVSG share a common semantic context in object-centric representation which making it well-suited for such modeling. At the heart of UNO is an extended slot attention mechanism [29] that decomposes visual feature maps into compact object and relation slots. These slots serve as modular building blocks and form a shared latent representation space across both tasks. To ensure temporal consistency, we introduce object temporal consistency learning, which enforces the alignment of object slots across frames without the need for explicit tracking. Furthermore, we propose a dynamic triplet prediction module that efficiently associates relation slots with subject–object pairs while reducing redundancy in the predicted triplets.
We validate UNO on two standard benchmarks: Action Genome [17] for DSGG task and PVSG [58] for PVSG task. Experimental results show that UNO consistently outperforms state-of-the-art (SOTA) methods in both accuracy and computational efficiency. To the best of our knowledge, UNO is among the early attempts to develop a unified, one-stage framework that jointly addresses box-level and pixel-level VidSGG within an object-centric paradigm.
@@
STTran [7] ICCV’21
APT [25] CVPR’22
TEMPURA [32] CVPR’23
PVSG [58] CVPR’23
OED [47] CVPR’24
MCL [34] AAAI’25
DIFFVSGG [3] CVPR’25
VISA [26] CVPR’25
@@ 2 Related Works
Video Scene Graph Generation (VidSGG). VidSGG is an extension of Scene Graph Generation [2] that analyses videos to identify objects and their relationships, representing this information as a structured graph to support high-level video understanding tasks [49, 41, 62, 36, 39, 55]. Researchers have explored how VidSGG can be leveraged on different granularities of video content, from coarse bounding boxes [17, 8, 27, 45, 7] to fine-grained panoptic masks [58, 50, 57], to represent dynamic interactions among objects with varying levels of precision. In fact, the literature has largely diverged into two directions:
@@
Dynamic Scene Graph Generation–DSGG [17, 8, 27, 45, 7] adopts the box-level approach to VidSGG that involves detection and tracking of object instances, capturing both spatial and temporal relationships to form graphs. In particular, Action Genome [17] supports DSGG with bounding boxes, relationship labels, and actions of human-object interactions. Various strategies have been proposed to predict objects and classify their pair-wise relationships [8, 27, 7]. Recently, OED [47] reformulates DSGG as a set prediction problem on object boxes, and leverages pair-wise features to represent each subject-object pair within the scene graph. We refer to this as box-level VidSGG.
@@
Video Panoptic Scene Graph Generation–PVSG [58, 50, 57] requires nodes in scene graphs to be grounded by precise, pixel-level segmentation masks to facilitate fine-grained scene understanding. In the PVSG benchmark [58], frames in a video are assigned with panoptic masks that provide pixel-level detail of object and background boundaries, where methods [58, 50] have had to capture both evolving object-background interactions and object-object interactions to create a cohesive, dynamic scene graph representation. We refer to this as pixel-level VidSGG.
@@
Interestingly, VidSGG strategies are often multi-stage pipelines [8, 27, 45, 7, 50, 57], while OED [47] recently pioneered a one-stage VidSGG approach for DSGG, eliminating the need for external tracking or multi-stage optimization. Although OED could not be directly adopted into PVSG, it highlights an important design consideration for stable learning and efficient end-to-end VidSGG modeling. Building on this insight, our research introduces a novel one-stage framework that unifies dynamic scene graph generation for both DSGG and PVSG in videos, setting it apart from prior approaches.
@@
Object-centric Representation Learning. Object-centric representation learning has been adopted to focus on entities [33, 42, 56, 36, 21] that are directly meaningful for study predictions. It is employed to uncover modular structures and independent mechanisms, such as objects and their relationships, from multi-object visual inputs [19, 38, 60]. Conventional models rely on object-/region-specific priors to facilitate reasoning and comprehension [33, 61, 42, 56]. Meanwhile, recent works have leveraged slot attention [38, 21, 29] to facilitate object-centric representations from raw scene features. In particular, slot attention has used GRU [6] and competitive attention mechanisms to bind to modular structures in the input [19, 29], potentially maintaining them through time for video understanding [38]. However, only a few studies have been done to consider an object-centric perspective for VidSGG besides straightforward visual detection, despite how it can consistently disentangle object semantics from general scene details. In this work, our research aims to extend their utility to capture objects’ modular structures and relationships for a unified VidSGG.
@@ 3.1 Preliminary and Motivation
Current VidSGG tasks. Given a video $\mathbf{V}=\langle I_{1},\ldots,I_{T}\rangle$ of $T$ frames, where each frame $I_{t}\in\mathbb{R}^{H_{\mathtt{in}}\times W_{\mathtt{in}}\times 3}$, VidSGG aims to produce a sequence of scene graphs $\mathbf{G}=\langle G_{1},\ldots,G_{T}\rangle$, where each $G_{t}$ represents the scene graph for frame $I_{t}$. Each $G_{t}$ consists of triplets and is defined as $G_{t}=\{\textit{subject},\textit{relation},\textit{object}\}$. The goal is to model the conditional probability $\mathbb{P}(\mathbf{G}\mid\mathbf{V})$. To simplify, existing methods [25] reformulate this as predicting detected objects and their pairwise relations.
$\mathbb{P}(\mathbf{G}\mid\mathbf{V})=\mathbb{P}(\mathbf{B},\mathbf{O},\mathbf{R}\mid\mathbf{V}).$ (1)
Here, $\mathbf{B}=\langle\mathbf{B}_{1},\ldots,\mathbf{B}_{T}\rangle$ is the set of bounding boxes for entire video $\mathbf{V}$ , where $\mathbf{B}_{t}=\{B_{1},\ldots,B_{M_{t}}\}$ is the set of bounding boxes of $M_{t}$ objects in the $t$-frame. Similarly, $\mathbf{O}=\langle\mathbf{O}_{1},\ldots,\mathbf{O}_{T}\rangle$ is the set of object labels for entire video $\mathbf{V}$ , where $\mathbf{O}_{t}=\{O_{1},\ldots,O_{M_{t}}\}$ is the set of object labels for the $t$-frame. $\mathbf{R}=\langle\mathbf{R}_{1},\ldots,\mathbf{R}_{T}\rangle$ is the set of relation for entire video $\mathbf{V}$ , where $\mathbf{R}_{t}=\{R_{1},\ldots,R_{L}\}$ is the set of relations for the $t$-frame. This approach is termed as coarse-grained box-level VidSGG task (DSGG [17]), and the conditional probability is factorized as follows:
@@
A prior study [58] extended this formulation to handle the fine-grained pixel-level VidSGG by replacing the bounding boxes $\mathbf{B}$ with mask tubes $\mathbf{M}$, which represent each object in the entire video. This leads to the formulation: $\mathbb{P}(\mathbf{G}\mid\mathbf{V})=\mathbb{P}(\mathbf{M},\mathbf{O},\mathbf{R}\mid\mathbf{V})$. We refer to this as the fine-grained pixel-level VidSGG task (PVSG [58]), and the conditional probability is factorized as follows.
@@
Unifying VidSGG tasks. Given the current formulations of DSGG in Eq.2 and PVSG in Eq.3, the primary distinction lies in their temporal granularity. DSGG operates at the frame level, where $\mathbf{O}_{t}$ denotes the set of objects detected within each individual frame, and thus does not enforce temporal consistency across frames. In contrast, PVSG leverages mask tubes, which inherently depend on maintaining object consistency throughout the video. A naïve approach to unifying VidSGG would result in a multi-stage pipeline of suboptimal performance and computational cost due to different modeling strategies in each task and stage. To address this challenge, in this work, we consider two VidSGG tasks through a unified perspective and propose a framework to directly model $\mathbb{P}\left(\mathbf{G}\mid\mathbf{V}\right)$ such that $\mathbf{G}=\{\mathbf{G_{\text{DSGG}}},\mathbf{G_{\text{PVSG}}}\}$ in an end-to-end training and inference manner. The primary challenge of our research lies in maintaining unified, spatio-temporal representations that align with video dynamics across both tasks, which we propose to address through an object-centric design, equipped with an object temporal consistency learning mechanism, ensuring a consistent and structured representation of objects and their relationships throughout the video sequence.
Design Principles. We highlight several key design principles of UNO. First, UNO is a one-stage, unified framework that minimizes task-specific modifications and multi-stage training/inference while maximizing parameter sharing via object-centric representation. Second, UNO serves as a novel baseline capable of handling diverse VidSGG tasks. By unifying VidSGG within a single framework, our approach reduces computational costs compared to multi-stage methods while maintaining competitive performance, establishing a strong foundation for current and future VidSGG research.

### v3

@@ Abstract
Video Scene Graph Generation (VidSGG) aims to represent dynamic visual content by detecting objects and modeling their temporal interactions as structured graphs. Prior studies typically target either coarse-grained box-level or fine-grained panoptic pixel-level VidSGG, often requiring task-specific architectures and multi-stage training pipelines. In this paper, we present UNO (UNified Object-centric VidSGG), a single-stage, unified framework that jointly addresses both tasks within an end-to-end architecture. UNO is designed to minimize task-specific modifications and maximize parameter sharing, enabling generalization across different levels of visual granularity. The core of UNO is an extended slot attention mechanism that decomposes visual features into object and relation slots. To ensure robust temporal modeling, we introduce object temporal consistency learning, which enforces consistent object representations across frames without relying on explicit tracking modules. Additionally, a dynamic triplet prediction module links relation slots to corresponding object pairs, capturing evolving interactions over time. We evaluate UNO on standard box-level and pixel-level VidSGG benchmarks. Results demonstrate that UNO not only achieves competitive performance across both tasks but also offers improved efficiency through a unified, object-centric design.
@@ 1 Introduction
Video Scene Graph Generation (VidSGG) aims to extract structured, dynamic representations from videos by modeling objects as nodes and their pairwise interactions as edges in spatio-temporal graphs. These structured representations offer both interpretability and compositionality, making VidSGG a critical component in various downstream tasks such as video understanding [ji2020action, av_scenegraph2024, yang2023pvsg, actionsg4ego2024], video reasoning [sc_actionreasoning2019, sganticipation2024] and robotic reasoning [shuran2020dynamic4robot].
@@
Current VidSGG research primarily follows two directions, distinguished by the level of visual granularity: box-level VidSGG (also referred to as Dynamic Scene Graph Generation or DSGG) [ji2020action, Wang_2024_CVPR, nag2023unbiased, li2022dynamic, teng2021target], and panoptic pixel-level VidSGG (also known as Panoptic Video Scene Graph Generation or PVSG) [yang2023pvsg, panopticsg2024tpami, psgg4D_2023]. The former focuses on coarse-grained object representations using bounding boxes and typically models relationships at the frame level. The latter provides fine-grained, pixel-level representations using panoptic segmentation masks, where object trajectories are treated as graph nodes and interactions—including object-object and object-background—are captured throughout the video. Importantly, PVSG emphasizes the temporal consistency of object identities across frames.
@@
In scenarios requiring multi-level scene understanding, a unified VidSGG model capable of handling both tasks is highly desirable. Such a model could flexibly adapt to diverse visual representations and support a wider range of applications without task-specific architectural redesign. However, achieving this unification is non-trivial due to the differing structural assumptions, temporal modeling requirements, and visual encoding strategies inherent to each task. Prior attempts have relied on multi-stage pipelines involving either box-level or pixel-level representation, followed by a tracking module [li2022dynamic, wang2023cross, zhang2023end, yang2023pvsg, mcl2025aaai], which introduce significant computational overhead and often result in suboptimal performance due to decoupled learning and limited parameter sharing. On the other hand, designing a unified, end-to-end solution encounters the core challenge of learning a semantically consistent spatio-temporal representation that generalizes across varying levels of granularity while remaining aligned with the underlying video dynamics.
@@
Beyond spatial granularity, VidSGG approaches also differ in their temporal-level representations, which can be broadly categorized into frame-level [cong2021spatial, li2022dynamic, nag2023unbiased, Wang_2024_CVPR] and tracklet-level [yang2023pvsg, mcl2025aaai] methods. Frame-level methods construct scene graphs independently at each frame, aligning naturally with box-level VidSGG, but often fall short in modeling long-term interactions and maintaining temporal consistency. In contrast, tracklet-level methods are more accurate because they link object instances across frames to capture temporal dynamics explicitly, a strategy more common in pixel-level VidSGG. This lack of temporal generality hinders their ability to capture coherent and continuous scene dynamics—especially in applications requiring both fine-grained relationship modeling and long-range temporal reasoning. A detailed comparison of UNO with prior works is presented in Tab. 1.
@@
STTran [cong2021spatial] ICCV’21
APT [li2022dynamic] CVPR’22
TEMPURA [nag2023unbiased] CVPR’23
PVSG [yang2023pvsg] CVPR’23
OED [Wang_2024_CVPR] CVPR’24
MCL [mcl2025aaai] AAAI’25
DIFFVSGG [DIFFVSGG] CVPR’25
VISA [visa2025] CVPR’25
@@
To address this challenge, we propose UNO, a unified, object-centric, single-stage VidSGG framework that effectively supports both box-level and pixel-level tasks. Fig. 1 illustrates the overall concept of UNO. Our central hypothesis is that despite their differences, DSGG and PVSG share a common semantic context in object-centric representation which making it well-suited for such modeling. At the heart of UNO is an extended slot attention mechanism [locatello2020object] that decomposes visual feature maps into compact object and relation slots. These slots serve as modular building blocks and form a shared latent representation space across both tasks. To ensure temporal consistency, we introduce object temporal consistency learning, which enforces the alignment of object slots across frames without the need for explicit tracking. Furthermore, we propose a dynamic triplet prediction module that efficiently associates relation slots with subject–object pairs while reducing redundancy in the predicted triplets.
We validate UNO on two standard benchmarks: Action Genome [ji2020action] for DSGG task and PVSG [yang2023pvsg] for PVSG task. Experimental results show that UNO consistently outperforms state-of-the-art (SOTA) methods in both accuracy and computational efficiency. While UNO builds on prior concepts such as slot attention and temporal contrastive consistency, our key contribution lies in recontextualizing and integrating them into a unified, one-stage framework specifically designed for VidSGG. To the best of our knowledge, UNO is among the first approaches to jointly address both box-level and pixel-level VidSGG within an object-centric paradigm.
@@ 2 Related Works
Video Scene Graph Generation (VidSGG). VidSGG is an extension of Scene Graph Generation [sggreview2023] that analyses videos to identify objects and their relationships, representing this information as a structured graph to support high-level video understanding tasks [av_scenegraph2024, actionsg4ego2024, sc_actionreasoning2019, sganticipation2024, sc_actionreg2021, shuran2020dynamic4robot]. Researchers have explored how VidSGG can be leveraged on different granularities of video content, from coarse bounding boxes [ji2020action, deng2019relation, lin2020gps, teng2021target, cong2021spatial] to fine-grained panoptic masks [yang2023pvsg, panopticsg2024tpami, psgg4D_2023], to represent dynamic interactions among objects with varying levels of precision. In fact, the literature has largely diverged into two directions:
@@
Dynamic Scene Graph Generation–DSGG [ji2020action, deng2019relation, lin2020gps, teng2021target, cong2021spatial] adopts the box-level approach to VidSGG that involves detection and tracking of object instances, capturing both spatial and temporal relationships to form graphs. In particular, Action Genome [ji2020action] supports DSGG with bounding boxes, relationship labels, and actions of human-object interactions. Various strategies have been proposed to predict objects and classify their pair-wise relationships [deng2019relation, lin2020gps, cong2021spatial]. Recently, OED [Wang_2024_CVPR] reformulates DSGG as a set prediction problem on object boxes, and leverages pair-wise features to represent each subject-object pair within the scene graph. We refer to this as box-level VidSGG.
@@
Video Panoptic Scene Graph Generation–PVSG [yang2023pvsg, panopticsg2024tpami, psgg4D_2023] requires nodes in scene graphs to be grounded by precise, pixel-level segmentation masks to facilitate fine-grained scene understanding. In the PVSG benchmark [yang2023pvsg], frames in a video are assigned with panoptic masks that provide pixel-level detail of object and background boundaries, where methods [yang2023pvsg, panopticsg2024tpami] have had to capture both evolving object-background interactions and object-object interactions to create a cohesive, dynamic scene graph representation. We refer to this as pixel-level VidSGG.
@@
Interestingly, VidSGG strategies are often multi-stage pipelines [deng2019relation, lin2020gps, teng2021target, cong2021spatial, panopticsg2024tpami, psgg4D_2023], while OED [Wang_2024_CVPR] recently pioneered a one-stage VidSGG approach for DSGG, eliminating the need for external tracking or multi-stage optimization. Although OED could not be directly adopted into PVSG, it highlights an important design consideration for stable learning and efficient end-to-end VidSGG modeling. Building on this insight, our research introduces a novel one-stage framework that unifies dynamic scene graph generation for both DSGG and PVSG in videos, setting it apart from prior approaches.
@@
Object-centric Representation Learning. Object-centric representation learning has been adopted to focus on entities [objectcentricprior4act2021, whatwhere2024, egoexoalign2023, sganticipation2024, actionslot2024] that are directly meaningful for study predictions. It is employed to uncover modular structures and independent mechanisms, such as objects and their relationships, from multi-object visual inputs [slotssm2024, rethinkimage2video2024, zhou2022slot]. Conventional models rely on object-/region-specific priors to facilitate reasoning and comprehension [objectcentricprior4act2021, viola2022, whatwhere2024, egoexoalign2023]. Meanwhile, recent works have leveraged slot attention [rethinkimage2video2024, actionslot2024, locatello2020object] to facilitate object-centric representations from raw scene features or for embodied and robotic features [slotvla, rethinkingoc]. In particular, slot attention has used GRU [chung2014empiricalgru] and competitive attention mechanisms to bind to modular structures in the input [slotssm2024, locatello2020object], potentially maintaining them through time for video understanding [rethinkimage2video2024]. However, only a few studies have been done to consider an object-centric perspective for VidSGG besides straightforward visual detection, despite how it can consistently disentangle object semantics from general scene details. In this work, our research aims to extend their utility to capture objects’ modular structures and relationships for a unified VidSGG.
@@ 3.1 Preliminary and Motivation
Current VidSGG tasks. Given a video $\mathbf{V}=\langle I_{1},\ldots,I_{T}\rangle$ of $T$ frames, where each frame $I_{t}\in\mathbb{R}^{H_{\mathtt{in}}\times W_{\mathtt{in}}\times 3}$, VidSGG aims to produce a sequence of scene graphs $\mathbf{G}=\langle G_{1},\ldots,G_{T}\rangle$, where each $G_{t}$ represents the scene graph for frame $I_{t}$. Each $G_{t}$ consists of triplets and is defined as $G_{t}=\{\textit{subject},\textit{relation},\textit{object}\}$. The goal is to model the conditional probability $\mathbb{P}(\mathbf{G}\mid\mathbf{V})$. To simplify, existing methods [li2022dynamic] reformulate this as predicting detected objects and their pairwise relations.
$\small\mathbb{P}(\mathbf{G}\mid\mathbf{V})=\mathbb{P}(\mathbf{B},\mathbf{O},\mathbf{R}\mid\mathbf{V}).$ (1)
Here, $\mathbf{B}=\langle\mathbf{B}_{1},\ldots,\mathbf{B}_{T}\rangle$ is the set of bounding boxes for entire video $\mathbf{V}$ , where $\mathbf{B}_{t}=\{B_{1},\ldots,B_{M_{t}}\}$ is the set of bounding boxes of $M_{t}$ objects in the $t$-frame. Similarly, $\mathbf{O}=\langle\mathbf{O}_{1},\ldots,\mathbf{O}_{T}\rangle$ is the set of object labels for entire video $\mathbf{V}$ , where $\mathbf{O}_{t}=\{O_{1},\ldots,O_{M_{t}}\}$ is the set of object labels for the $t$-frame. $\mathbf{R}=\langle\mathbf{R}_{1},\ldots,\mathbf{R}_{T}\rangle$ is the set of relation for entire video $\mathbf{V}$ , where $\mathbf{R}_{t}=\{R_{1},\ldots,R_{L}\}$ is the set of relations for the $t$-frame. This approach is termed as coarse-grained box-level VidSGG task (DSGG [ji2020action]), and the conditional probability is factorized as follows:
@@
A prior study [yang2023pvsg] extended this formulation to handle the fine-grained pixel-level VidSGG by replacing the bounding boxes $\mathbf{B}$ with mask tubes $\mathbf{M}$, which represent each object in the entire video. This leads to the formulation: $\mathbb{P}(\mathbf{G}\mid\mathbf{V})=\mathbb{P}(\mathbf{M},\mathbf{O},\mathbf{R}\mid\mathbf{V})$. We refer to this as the fine-grained pixel-level VidSGG task (PVSG [yang2023pvsg]), and the conditional probability is factorized as follows.
@@
Unifying VidSGG tasks. Given the current formulations of DSGG in Eq.2 and PVSG in Eq.3, the primary distinction lies in their temporal granularity. DSGG operates at the frame level, where $\mathbf{O}_{t}$ denotes the set of objects detected within each individual frame, and thus does not enforce temporal consistency across frames. In contrast, PVSG leverages mask tubes, which inherently depend on maintaining object consistency throughout the video. A naïve approach to unifying VidSGG would result in a multi-stage pipeline of suboptimal performance and computational cost due to different modeling strategies in each task and stage. To address this challenge, in this work, we consider two VidSGG tasks through a unified perspective and propose a framework to directly model $\mathbb{P}\left(\mathbf{G}\mid\mathbf{V}\right)$ such that $\mathbf{G}=\{\mathbf{G_{\text{DSGG}}}\text{ or }\mathbf{G_{\text{PVSG}}}\}$ in an end-to-end training and inference manner. The primary challenge of our research lies in maintaining unified, spatio-temporal representations that align with video dynamics across both tasks, which we propose to address through an object-centric design, equipped with an object temporal consistency learning mechanism, ensuring a consistent and structured representation of objects and their relationships throughout the video sequence.
Design Principles. UNO follows three key principles. First, it is a one-stage unified framework that minimizes task-specific modifications and multi-stage processing while maximizing parameter sharing through object-centric representation. Second, it establishes a strong baseline for diverse VidSGG tasks by reducing computational cost compared to multi-stage methods without sacrificing performance. Finally, instead of fusing bounding boxes and masks at the output level, UNO employs a unified object-centric representation using Slot Attention, where object and relation slots serve as shared latent features and feed into task-specific heads for bounding box or panoptic mask prediction.